# Security of Language Models for Code: A Systematic Literature Review

### v1

@@ 4.2.1. Overview of Backdoor Attacks against CodeLMs
Cotroneo et al. (Cotroneo et al., 2024) introduce a targeted data poisoning attack method to assess the security of Natural Language to Code (NL-to-Code) generators by injecting software vulnerabilities into fine-tuned model data. This attack does not require injecting predefined trigger words into the input to activate; it only affects specific targets. Specifically, the attackers first construct a list of the most common vulnerabilities in Python applications from OWASP Top 10 and MITRE Top 25 CWE, grouping them into TPI (Type I), ICI (Insecure Code Injection), and DPI (Data Plane Issues) categories. Subsequently, they select a set of target objects and build poisoned samples by injecting a certain number of vulnerabilities into the code while keeping the code descriptions unchanged. They evaluate the impact of model architecture, poisoning rate, and vulnerability categories on neural machine translation (NMT) models using CodeLMs like Seq2Seq, CodeBERT, and CodeT5+ (Wang et al., 2023c). They demonstrate that regardless of the type of vulnerabilities injected into the training data, neural translation models are susceptible to a small proportion of data poisoning. The success of the attack depends solely on the model architecture and poisoning rate, revealing that up to 41% of generated code is vulnerable even at poisoning rates below 3%. Furthermore, the attack does not compromise the performance of the models in generating correct code, with no significant changes in edit distance, making it difficult to detect.
Zhou et al. (Yang et al., 2024c) introduce a covert backdoor attack method named AFRAIDOOR targeted at CodeLMs. AFRAIDOOR employs adversarial perturbations, identifier renaming, and adaptive triggers to ensure the stealthiness of the attack. Specifically, it initially trains a Seq2Seq model on a clean dataset. Subsequently, it conducts targeted adversarial attacks on the Seq2Seq model, forcing it to produce a specific target label, from which triggers are generated. During the attack, AFRAIDOOR replaces all local identifiers with a special token ([UNK]). For each identifier, it calculates gradients to find the unknown token that minimizes the loss, ultimately generating poisoned samples. AFRAIDOOR adaptively generates different triggers based on the data, making it more covert compared to fixed trigger and syntax-trigger backdoor attack methods and less prone to detection by defensive methods. Their experimental results demonstrate that 85% of adaptive triggers generated by AFRAIDOOR bypass detection using spectrum methods (Tran et al., 2018).
Aghakhani et al. (Aghakhani et al., 2024) propose two novel attacks, COVERT and TROJANPUZZLE. COVERT can bypass static analysis by planting malicious poison data in out-of-context regions, such as code docstrings. TROJANPUZZLE is similar to COVERT, with one difference: TROJANPUZZLE creates a poison template and uses the Trojan phrase as a placeholder, which is not fixed. Experiments demonstrate that COVERT yields result comparable to the SIMPLE attack (Schuster et al., 2021) that employs explicit poison code. TROJANPUZZLE demonstrates lower success rates compared to SIMPLE and COVERT, but is robust against signature-based dataset-cleansing methods (Schuster et al., 2021).
4.2.3. Model Poisoning Attack.
@@
Li et al. (Li et al., 2023b) propose a task-agnostic backdoor attack method targeting pre-trained CodeLMs. This attack constructs a trigger set containing both code and natural language elements and employs two strategies, poisoning sequence-to-sequence learning (Seq2Seq learning) and token representation learning, to poison the model during pre-training, enabling multi-objective attacks on code generation and understanding tasks. They chose to insert low-frequency words “cl” and “tp” as triggers in the natural language part of the data and insert a piece of dead code as triggers in the code part of the data. During the deployment phase, the implanted backdoor in the target model can be activated using the designed triggers, facilitating targeted attacks. They evaluate the attacks on two code understanding tasks and three code generation tasks across seven datasets. The experiments demonstrate that the attacks can effectively and covertly compromise downstream code-related tasks.
Different from the aforementioned study, Zhang et al. (Zhang et al., 2023d) investigate backdoor vulnerabilities in natural training CodeLMs used for binary code analysis. They design trigger inversion techniques for trigger generation and inject those triggers to model, then invent PELICAN as the prototype of the attack method. PELICAN initially constructs an instruction dictionary and uses gradient descent to search for instructions that may lead to misclassifications as triggers. Subsequently, it employs a randomized micro-execution technique to extract program state changes for a given binary code. Finally, it utilizes a solution-based approach to generate binary code that satisfies trigger injection requirements while preserving semantics. Importantly, the backdoor vulnerabilities in natural training models are not injected by attackers but rather stem from defects in the dataset or training process. They evaluate PELICAN on five binary code analysis tasks and 15 models. Experimental results demonstrate that PELICAN can effectively induce misclassifications in all evaluated models, achieving an attack success rate of 86.09% with only three trigger instructions. Compared to opaque predicates (Junod et al., 2015), which use non-transparent predicates as triggers, PELICAN reduces runtime by 204.23% while increasing the attack success rate by 93.01%. Additionally, 94.14% of the triggers injected by PELICAN can evade detection, whereas triggers injected through opaque predicates are all detectable.
4.2.4. Summary of Existing Backdoor Attacks
@@
BADCODE (Sun et al., 2023a) 2023 ACL Data poisoning CodeBERT CodeT5 Code search
@@
CodePoisoner (Li et al., 2024b) 2024 TOSEM Data poisoning LSTM TextCNN Transformer CodeBERT Defect detection Clone detection Code repair
AFRAIDOOR (Yang et al., 2024c) 2024 TSE Data poisoning CodeBERT CodeT5 PLBART Code summarization
@@
Li et al. (Li et al., 2023b) 2023 ACL Model poisoning PLBART CodeT5 Defect detection Clone prediction Code2Code translation Text2Code translation Code refine
@@
PELICAN (Zhang et al., 2023d) 2023 USENIX Security Model poisoning BiRNN-func XDA-func XDA-cell StateFormer EKLAVYA EKLAVYA++ in-nomine in-nomine++ S2V, S2V++ Trex SAFE, SAFE++ S2V-B, S2V-B++ Binary code analysis
@@
Table 3 summarizes the existing backdoor attacks on CodeLMs. Both data poisoning attacks and model poisoning attacks have demonstrated their effectiveness on various CodeLMs, including non-pre-trained and pre-trained models. Additionally, these attacks have proven effective across multiple code-related tasks, including code understanding and code generation tasks. Data poisoning is the most extensively studied method for backdoor attacks targeting CodeLMs. Among these, attacks based on fixed or syntactically invalid dead code snippets as triggers are the most common (Ramakrishnan and Albarghouthi, 2022; Wan et al., 2022; Qi et al., 2023; Li et al., 2024b). However, dead code triggers, typically composed of one or more lines of code, tend to have low stealth and can be easily detected by users, developers, or static analysis tools. In addition to research on dead code triggers, there are studies focused on the stealthiness of backdoor triggers and the naturalness of code after trigger implantation. These techniques often leverage code frequency, context, or adversarial perturbations and usually execute the attack by replacing identifiers. While these methods can enhance stealth, they generally have lower success rates compared to fixed dead code triggers and are often task-specific. Backdoor attacks based on model poisoning have gained attention since 2023 (Qi et al., 2023; Li et al., 2023b). Compared to data poisoning, model poisoning allows attackers to manipulate the training process and release backdoored pre-trained models on open-source platforms, posing a greater threat than data poisoning. However, current model poisoning attacks also implant fixed or syntactically invalid dead code snippets as triggers, facing similar challenges of insufficient stealth. In addition to intentional backdoor implantations, (Zhang et al., 2023d) has also explored the backdoor security of naturally trained models, revealing that even naturally trained models can have exploitable backdoors.
Sun et al. (Sun et al., 2024b) propose a novel attack method called Feedback Data Injection (FDI), which targets neural code generation systems by exploiting their feedback mechanisms. This attack method leverages user feedback to inject malicious code snippets into the system, leading to vulnerabilities such as backdoor attacks and prompt injection attacks. The FDI method allows an attacker to iteratively refine their malicious samples, increasing the attack’s success over time.
Yi et al. (Yi et al., 2024) propose a novel backdoor attack method called LateBA, which targets deep bug search models by injecting triggers into infrequently executed code. This approach improves the stealth of backdoor attacks by focusing on rarely executed code segments, which are less likely to be detected by dynamic analysis. LateBA employs a two-step process: first, it identifies infrequent execution code ranges through a path search tool like American Fuzzy Lop (AFL); second, it finds the optimal locations within these ranges for injecting the backdoor triggers. The method leverages semantic feature comparisons to ensure that the injected triggers do not interfere with the overall execution logic, thus maintaining high stealth and effectiveness. Experiments demonstrate that LateBA achieves a significant reduction in trigger detection and program crashes, while maintaining a high attack success rate across different optimization levels of binary files.
4.3. Adversarial Attacks on CodeLMs.
4.3.1. Overview of Adversarial Attacks against CodeLMs.
@@
4.3.2. White-box Attacks on CodeLMs.
@@
Bielik et al. (Bielik and Vechev, 2020) propose a new approach to improve the robustness while preserving high accuracy for CodeLMs in the task of type inference for two dynamic type languages, JavaScript and TypeScript. They allow the model to abstain from predicting parts of the samples, abstracting parts of the program to reduce complexity, and propose a strategy of training multiple models in parallel. Firstly, they perform semantic-preserving modifications and label-preserving modifications on code, which include renaming variables and altering program structures. Then, they generate adversarial samples using an iterative method and evaluate them on five neural model architectures, including LSTM, DeepTyper, GCN, etc. Experimental results demonstrate that their approach achieves an accuracy of 87%, while increasing the model’s robustness from 52% to 67%.
Srikant et al. (Srikant et al., 2021) propose an attack method that confuses transformations on code to generate adversarial perturbations, disrupting model judgments while preserving program semantics. They introduce a set of first-order optimization algorithms and a general formula for adversarial attacks, transforming adversarial attack generation into a constrained combinatorial optimization problem. Specifically, they first use first-order optimization algorithms to determine the location of perturbations in the program and specific perturbations, including replacing existing tokens or inserting new tokens. Simultaneously, they use the general formula for adversarial attacks on program obfuscation transformations for any programming language. Then, they propose a random smoothing algorithm to enhance optimization performance. They evaluate the attack on Java and Python function name prediction tasks using Seq2Seq and Code2Seq models. Results indicate that the attack success rate is 1.5 times higher than Ramakrishnan et al.’s (Henkel et al., 2022) attack generation algorithm. They also demonstrate that the proposed attack can be used for adversarial training to improve model adversarial robustness.
Zhang et al. (Zhang et al., 2022a) introduce an optimization-based adversarial attack technique named CARROTA, specifically tailored for CodeLMs based on gradient-guided approaches. CARROTA offers transformations at both token and statement levels, iteratively seeking adversarial examples capable of misleading the model into incorrect classifications. Differing from MHM, which perturbs only at the identifier level, CARROTA supports any grammar-compliant synonym transformations. Additionally, CARROTA combines gradient information to guide the adversarial sample search process with reduced computational cost. They also propose a metric called CARROTM for estimating robustness boundaries. They also present an adversarial training method, CARROTT, to periodically generate adversarial examples for augmenting the training set, integrating these techniques into a comprehensive framework named CARROT. CARROT enables systematic robustness assessment, measurement, and enhancement for CodeLMs. Extensive evaluations are conducted on models (e.g., GRU, LSTM, CodeBERT) across tasks like code functionality classification, code clone detection, and code defect prediction. Results indicate CARROTA’s efficiency in generating adversarial examples, resulting in an average performance decrease of 87.2% at the token level, surpassing MHM’s 75.5%. CARROTM provides more accurate robustness estimates under stricter constraints, obtaining significantly stricter robustness boundaries compared to simple random baseline methods and MHM. CARROTT enhances the robustness of CodeLMs, achieving an average improvement of 5.3 times over the original models, outperforming MHM by 1.7 times.
@@
Chen et al. (Chen et al., 2022) propose a novel method of generating adversarial examples in the white-box setting to ensure that the perturbed code retains the original semantics. First, they parse the code into a token sequence, calculate the contribution value of each token to the classification result using the Jacobian matrix, and determine the importance order of tokens based on their contribution values. Secondly, they select appropriate code transformations to apply to the perturbation positions of important tokens and choose the best transformation based on the model’s prediction results. The experimental results show that in the code summarization task, their attack method achieves an average improvement of 8% in attack success rate compared to alternating optimization with randomized smoothing (Srikant et al., 2021). Additionally, adversarial training using the adversarial examples generated by their method can reduce the attack success rates by an average of 21%.
@@
Capozzi et al. (Capozzi et al., 2023) investigate the resilience of binary similarity models against adversarial attacks, specifically focusing on white-box scenarios. The study employs a gradient-guided strategy, repurposing a method typically used in image classifier attacks, to insert instructions into binary code while maintaining semantic integrity. This approach is implemented by iteratively minimizing the model’s loss function while ensuring the perturbations remain small and semantically valid. The attack is tested against three state-of-the-art binary similarity solutions: Gemini (Xu et al., 2017), GMN (Li et al., 2019), and SAFE (Massarelli et al., 2021). The results reveal that these models are susceptible to both targeted and untargeted attacks from both black-box and white-box adversaries.
4.3.3. Black-box Attacks on CodeLMs.
@@
Matyukhina et al. (Matyukhina et al., 2019) introduce an author imitation attack to explore the potential of existing authorship attribution techniques to be deceived. The flow of the attack consists of three steps. Firstly, they identify a victim and retrieve samples of their source code. Once the samples are collected, the second step involves analyzing them and applying transformations to identify the victim’s coding style. Finally, they use these transformations to modify the original code, producing a semantically equivalent version that mimics the victim’s coding style. The experiments demonstrate that they can successfully launch an imitation attack on all users in Google Code Jam and the GitHub dataset.
Springer et al. (Springer et al., 2020) discover that the norm of the highest-frequency token embeddings learned by a model increases with the frequency of the tokens. They utilize this relationship to develop a gradient-free variable substitution adversarial attack method called STRATA for generating adversarial examples on CodeLMs. STRATA utilizes three strategies for selecting high-impact tokens, including aggregating all tokens, selecting tokens from the top N highest norm embedding vectors, and selecting tokens with the highest frequency. Additionally, STRATA presents three token substitution strategies, including single substitution, five different substitutions, and five identical substitutions. Experiments demonstrate that STRATA can effectively launch adversarial attacks under these strategies. Moreover, STRATA outperforms the gradient-based adversarial attack method proposed by Ramakrishnan et al. (Henkel et al., 2022).
@@
Liu et al. (Liu et al., 2021) propose an effective source code authorship disguise method, called SCAD. SCAD is more practical than the work by Quiring et al. (Quiring et al., 2019) which requires knowledge of the output probabilities or internal details of the target CodeLMs for source code authorship identification. Specifically, SCAD first trains a substitute model and develops a set of semantically equivalent transformations. Then, based on these transformations, SCAD makes slight lexical and syntactic feature changes to the original code to achieve disguise. Their experimental results show that on a real dataset containing 1,600 programmers, SCAD induces error rates of over 30% in CodeLMs for author identification based on Random Forest Classifiers and RNN classifiers. They point out the necessity of developing more robust stylistic features and identification CodeLMs to mitigate authorship disguise attacks.
@@
Ferretti et al. (Ferretti and Saletta, 2021) propose an evolutionary approach to assess the robustness of software vulnerability detection models. The fundamental idea of the method is to explore the potential feature representation space using the Grammatical Evolution (GE) algorithm (O’Neill et al., 2014). Initially, the method evolves code snippets by using the classifier’s output as the fitness function. Then, it examines how well the selected evolved individuals can effectively alter the classification of any instance in the labeled dataset. Finally, the method generates adversarial samples by injecting semantics while preserving the syntax. This method leverages evolutionary pressure rather than random sampling, thus effectively exploring the behavior of classifiers and identifying features that could potentially mislead the model. Their experiments indicate that code snippets generated by the GE algorithm can significantly guide the network to classify modified instances in a manner opposite to the true labels with high probability, reducing the precision/recall area under curve (P/R-AUC index) of the neural classifier from 0.42 to 0.06.

### v2

@@ 4.2.2. Data Poisoning Attacks
To avoid detection by dynamic analysis or context-based detectors, Yi et al. (Yi et al., 2024) propose a backdoor attack method called LateBA, which targets deep bug search models by injecting triggers into infrequently executed code. This approach improves the stealth of backdoor attacks by focusing on rarely executed code segments, which are less likely to be detected by dynamic analysis. LateBA employs a two-step process: first, it identifies infrequent execution code ranges through a path search tool like American Fuzzy Lop (AFL); second, it finds the optimal locations within these ranges for injecting the backdoor triggers. The method leverages semantic feature comparisons to ensure that the injected triggers do not interfere with the overall execution logic, thus maintaining high stealth and effectiveness. Experiments demonstrate that LateBA achieves a significant reduction in trigger detection and program crashes, while maintaining a high attack success rate across different optimization levels of binary files.
To explore the impact of user feedback mechanisms on the security of neural code generation, Sun et al. (Sun et al., 2024a) propose an attack method called Feedback Data Injection (FDI). FDI attacks by exploiting the feedback mechanisms of neural code generation systems. Specifically, FDI leverages user feedback to inject malicious code snippets into the system, leading to vulnerabilities such as backdoor attacks and prompt injection attacks. The FDI method allows an attacker to iteratively refine their malicious samples, increasing the attack’s success over time.
Zhang et al. (Zhang et al., 2023e) propose a backdoor-based SQL injection framework, TrojanSQL, specifically tailored for text-to-SQL systems. Targeting natural language interface to database, TrojanSQL implements two distinct SQL injection techniques: a boolean-based injection, which inserts an always-true OR-condition in the WHERE clause to bypass original query constraints, and a union-based injection, which appends union queries to exfiltrate private information. TrojanSQL employs a sketch-based insertion process to maintain syntactical correctness and ensure full executability of malicious SQL statements, while dynamically constructing prompts and payloads based on user queries and the database schema to maximize stealth and efficacy. Experimental results indicate that both finetuning-based and LLM-based parsers require only a small number of poisoned samples to achieve high attack success rates.
4.2.3. Model Poisoning Attacks
@@
Furthermore, Li et al. (Li et al., 2023c) propose a task-agnostic backdoor attack method targeting pre-trained CodeLMs. This attack constructs a trigger set containing both code and natural language elements and employs two strategies, poisoning sequence-to-sequence learning (Seq2Seq learning) and token representation learning, to poison the model during pre-training, enabling multi-objective attacks on code generation and understanding tasks. They chose to insert low-frequency words “cl” and “tp” as triggers in the natural language part of the data and insert a piece of dead code as triggers in the code part of the data. During the deployment phase, the implanted backdoor in the target model can be activated using the designed triggers, facilitating targeted attacks. They evaluate the attacks on two code understanding tasks and three code generation tasks across seven datasets. The experiments demonstrate that the attacks can effectively and covertly compromise downstream code-related tasks.
Different from the aforementioned study, Zhang et al. (Zhang et al., 2023b) investigate backdoor vulnerabilities in natural training CodeLMs used for binary code analysis. They design trigger inversion techniques for trigger generation and inject those triggers to model, then invent PELICAN as the prototype of the attack method. PELICAN initially constructs an instruction dictionary and uses gradient descent to search for instructions that may lead to misclassifications as triggers. Subsequently, it employs a randomized micro-execution technique to extract program state changes for a given binary code. Finally, it utilizes a solution-based approach to generate binary code that satisfies trigger injection requirements while preserving semantics. Importantly, the backdoor vulnerabilities in natural training models are not injected by attackers but rather stem from defects in the dataset or training process. They evaluate PELICAN on five binary code analysis tasks and 15 models. Experimental results demonstrate that PELICAN can effectively induce misclassifications in all evaluated models, achieving an attack success rate of 86.09% with only three trigger instructions. Compared to opaque predicates (Junod et al., 2015), which use non-transparent predicates as triggers, PELICAN reduces runtime by 204.23% while increasing the attack success rate by 93.01%. Additionally, 94.14% of the triggers injected by PELICAN can evade detection, whereas triggers injected through opaque predicates are all detectable.
4.2.4. Summary of Backdoor Attacks
@@
BadCode (Sun et al., 2023) 2023 ACL Data poisoning CodeBERT CodeT5 Code search
@@
TrojanSQL (Zhang et al., 2023e) 2023 EMNLP Data poisoning DuoRAT, LGESQL ISESQL, Proton T5-Large, T5-3B GPT-3 Text-to-SQL
@@
CodePoisoner (Li et al., 2024b) 2024 TOSEM Data poisoning LSTM, TextCNN Transformer CodeBERT Defect detection Clone detection Code repair
AFRAIDOOR (Yang et al., 2024c) 2024 TSE Data poisoning CodeBERT CodeT5, PLBART Code summarization
@@
FDI (Sun et al., 2024a) 2024 ISSTA Data poisoning GPT-3 Code generation
LateBA (Yi et al., 2024) 2024 Internetware Data poisoning Transformer Deep bug search
@@
Li et al. (Li et al., 2023c) 2023 ACL Model poisoning PLBART CodeT5 Defect detection Clone prediction Code2Code translation Text2Code translation Code refine
@@
PELICAN (Zhang et al., 2023b) 2023 USENIX Security Model poisoning BiRNN-func, StateFormer XDA-func, XDA-cell EKLAVYA, EKLAVYA++ in-nomine, in-nomine++ S2V, S2V++, Trex SAFE, SAFE++ S2V-B, S2V-B++ Binary code analysis
@@
Table 4 summarizes the existing backdoor attacks on CodeLMs. Both data poisoning attacks and model poisoning attacks have demonstrated their effectiveness on various CodeLMs, including non-pre-trained and pre-trained models. Additionally, these attacks have proven effective across multiple code-related tasks, including code understanding and code generation tasks. Data poisoning is the most extensively studied method for backdoor attacks targeting CodeLMs. Among these, attacks based on fixed or syntactically invalid dead code snippets as triggers are the most common (Ramakrishnan and Albarghouthi, 2022; Wan et al., 2022; Qi et al., 2023; Li et al., 2024b). However, dead code triggers, typically composed of one or more lines of code, tend to have low stealth and can be easily detected by users, developers, or static analysis tools. In addition to research on dead code triggers, there are studies focused on the stealthiness of backdoor triggers and the naturalness of code after trigger implantation. These techniques often leverage code frequency, context, or adversarial perturbations and usually execute the attack by replacing identifiers. While these methods can enhance stealth, they generally have lower success rates compared to fixed dead code triggers and are often task-specific. Backdoor attacks based on model poisoning have gained attention since 2023 (Qi et al., 2023; Li et al., 2023c). Compared to data poisoning, model poisoning allows attackers to manipulate the training process and release backdoored pre-trained models on open-source platforms, posing a greater threat than data poisoning. However, current model poisoning attacks also implant fixed or syntactically invalid dead code snippets as triggers, facing similar challenges of insufficient stealth. In addition to intentional backdoor implantations, (Zhang et al., 2023b) has also explored the backdoor security of naturally trained models, revealing that even naturally trained models can have exploitable backdoors.
4.3. Adversarial Attacks on CodeLMs
4.3.1. Overview of Adversarial Attacks against CodeLMs
@@
4.3.2. White-box Attacks on CodeLMs
@@
To ensure that the perturbed code retains the original semantics, Chen et al. (Chen et al., 2022) propose a method of generating adversarial examples in the white-box setting. First, they parse the code into a token sequence, calculate the contribution value of each token to the classification result using the Jacobian matrix, and determine the importance order of tokens based on their contribution values. Secondly, they select appropriate code transformations to apply to the perturbation positions of important tokens and choose the best transformation based on the model’s prediction results. The experimental results show that in the code summarization task, their attack method achieves an average improvement of 8% in attack success rate compared to alternating optimization with randomized smoothing (Srikant et al., 2021). Additionally, adversarial training using the adversarial examples generated by their method can reduce the attack success rates by an average of 21%.
Bielik et al. (Bielik and Vechev, 2020) propose an approach to improve the robustness while preserving high accuracy for CodeLMs in the task of type inference for two dynamic type languages, JavaScript and TypeScript. They allow the model to abstain from predicting parts of the samples, abstracting parts of the program to reduce complexity, and propose a strategy of training multiple models in parallel. Firstly, they perform semantic-preserving modifications and label-preserving modifications on code, which include renaming variables and altering program structures. Then, they generate adversarial samples using an iterative method and evaluate them on five neural model architectures, including LSTM, DeepTyper, GCN, etc. Experimental results demonstrate that their approach achieves an accuracy of 87%, while increasing the model’s robustness from 52% to 67%.
To disrupt the model’s judgment without compromising the program’s semantics, Srikant et al. (Srikant et al., 2021) propose an attack method that generates adversarial perturbations through obfuscation transformations on the code. They introduce a set of first-order optimization algorithms and a general formula for adversarial attacks, transforming adversarial attack generation into a constrained combinatorial optimization problem. Specifically, they first use first-order optimization algorithms to determine the location of perturbations in the program and specific perturbations, including replacing existing tokens or inserting new tokens. Simultaneously, they use the general formula for adversarial attacks on program obfuscation transformations for any programming language. Then, they propose a random smoothing algorithm to enhance optimization performance. They evaluate the attack on Java and Python function name prediction tasks using Seq2Seq and Code2Seq models. Results indicate that the attack success rate is 1.5 times higher than Henkel et al.’s (Henkel et al., 2022) attack generation algorithm. They also demonstrate that the proposed attack can be used for adversarial training to improve model adversarial robustness.
@@
To investigate the resilience of binary similarity models against adversarial attacks, Capozzi et al. (Capozzi et al., 2023) propose a white-box attack method. This approach employs a gradient-guided strategy, repurposing a method typically used in image classifier attacks, to insert instructions into binary code while maintaining semantic integrity. This approach is implemented by iteratively minimizing the model’s loss function while ensuring the perturbations remain small and semantically valid. The attack is tested against three state-of-the-art binary similarity solutions: Gemini (Xu et al., 2017), GMN (Li et al., 2019), and SAFE (Massarelli et al., 2021). The results reveal that these models are susceptible to both targeted and untargeted attacks from both black-box and white-box adversaries.
4.3.3. Black-box Attacks on CodeLMs
@@
Matyukhina et al. (Matyukhina et al., 2019) also introduce an author imitation attack to explore the potential of existing authorship attribution techniques to be deceived. The flow of the attack consists of three steps. Firstly, they identify a victim and retrieve samples of their source code. Once the samples are collected, the second step involves analyzing them and applying transformations to identify the victim’s coding style. Finally, they use these transformations to modify the original code, producing a semantically equivalent version that mimics the victim’s coding style. The experiments demonstrate that they can successfully launch an imitation attack on all users in Google Code Jam and the GitHub dataset.
To reduce the reliance on the output probabilities or internal details of the target CodeLMs, Liu et al. (Liu et al., 2021) propose a source code authorship disguise method, called SCAD. Specifically, SCAD first trains a substitute model and develops a set of semantically equivalent transformations. Then, based on these transformations, SCAD makes slight lexical and syntactic feature changes to the original code to achieve disguise. Their experimental results show that on a real dataset containing 1,600 programmers, SCAD induces error rates of over 30% in CodeLMs for author identification based on Random Forest Classifiers and RNN classifiers. They point out the necessity of developing more robust stylistic features and identification CodeLMs to mitigate authorship disguise attacks.
Furthermore, Gao et al. (Gao et al., 2023) propose an adversarial attack named DaK, specifically designed for CodeLMs. DaK creates discrete adversarial examples by preserving semantic code input through program transformations. DaK comprises three key components: Destroyer, Finder, and Merger. Destroyer alters the functionality of input code to weaken the features on which the model relies for predictions, using only semantically preserving transformations. Finder identifies a set of code predicted as the target label, computes critical features within this set, and selects features deemed the strongest and most concentrated in influencing the model’s prediction. Finally, Merger injects the selected critical features into the target program, generating discrete adversarial examples. Experiments on Code2Vec, GGNN, and CodeBERT demonstrate that DaK outperforms Imitator (Quiring et al., 2019) and DAMP (Yefet et al., 2020) on attacking undefended or unreinforced CodeLMs.
Springer et al. (Springer et al., 2020) discover that the norm of the highest-frequency token embeddings learned by a model increases with the frequency of the tokens. They utilize this relationship to develop a gradient-free variable substitution adversarial attack method called STRATA for generating adversarial examples on CodeLMs. STRATA utilizes three strategies for selecting high-impact tokens, including aggregating all tokens, selecting tokens from the top N highest norm embedding vectors, and selecting tokens with the highest frequency. Additionally, STRATA presents three token substitution strategies, including single substitution, five different substitutions, and five identical substitutions. Experiments demonstrate that STRATA can effectively launch adversarial attacks under these strategies. Moreover, STRATA outperforms the gradient-based adversarial attack method proposed by Henkel et al. (Henkel et al., 2022).
@@
To further enhance the naturalness of adversarial examples, Yang et al. (Yang et al., 2022) introduce ALERT, a black-box adversarial attack. ALERT generates adversarial examples by modifying code variables, inducing the victim model to produce incorrect outputs by altering the input. Specifically, ALERT first uses CodeBERT and GraphCodeBERT pre-trained models to generate candidate alternative words for code variables. It then ranks these candidate alternative words based on cosine similarity and uses a greedy algorithm or a genetic algorithm to select adversarial examples from the candidate alternative words. Therefore, ALERT generates adversarial examples that retain the natural semantics and operational semantics of the code. Experiments demonstrate that ALERT outperforms MHM in attacking CodeBERT and GraphCodeBERT in vulnerability prediction, clone detection, and source code authorship attribution tasks.
Furthermore, Yu et al. (Yu et al., 2023) introduce a black-box adversarial attack framework, AdVulCode, specifically designed for generating code adversarial examples in the context of vulnerability detection tasks based on CodeLMs. Unlike ALERT (Yang et al., 2022) and MHM (Zhang et al., 2020a), which generate adversarial examples by perturbing identifiers, AdVulCode utilizes equivalent transformation rules that do not alter the semantic meaning of the source code. It targets susceptible code segments for transformation, thus reducing the overall perturbation level of the source code. Specifically, AdVulCode employs equivalent transformations to generate candidate code statements and introduces an improved Monte Carlo search tree to guide the selection of candidate statements and generate adversarial examples. In comparison to ALERT, AdVulCode can handle models with identifier name mappings. Experimental results demonstrate that AdVulCode outperforms ALERT in adversarial attacks for vulnerability detection tasks.
Nguyen et al. (Nguyen et al., 2023) propose a black-box adversarial attack framework called GraphCodeAttack. To assess the robustness of CodeLMs, GraphCodeAttack automatically discovers important code patterns influencing model decisions to perturb the code structure of input models. Specifically, GraphCodeAttack identifies discriminative abstract syntax tree (AST) patterns that influence model decisions from source code and target model outputs. It then selects suitable patterns to insert dead code into the source code, creating adversarial examples. Experimental results show that GraphCodeAttack achieves an average attack success rate of 0.841 on author attribution, code vulnerability detection, and code clone detection tasks for CodeBERT and GraphCodeBERT CodeLMs, surpassing ALERT and CARROTA adversarial attack methods.
Zhang et al. (Zhang et al., 2023a) propose RNNS, a black-box adversarial attack method designed to assess the robustness of pre-trained CodeLMs by generating adversarial test data based on model uncertainty and output variation. Similar to MHM (Zhang et al., 2020a) and ALERT (Yang et al., 2022), RNNS generates adversarial examples by renaming local variables in the source code. In contrast to MHM and ALERT, RNNS collects variable names from real code to generate statistically characteristic adversarial examples while pursuing a high attack success rate. Specifically, RNNS first constructs a search space for variable names based on semantic similarity using real code data. Then, it utilizes a Representation Nearest Neighbor Search method to find adversarial examples from the namespace guided by model output changes. Failed attack samples guide the next round of attacks in RNNS. Experimental results indicate that, compared to MHM and ALERT, RNNS generates adversarial instances with smaller perturbations and achieves superior attack effectiveness on CodeBERT, GraphCodeBERT, and CodeT5 across six code tasks in three programming languages.
@@
Zhang et al. (Zhang et al., 2022a) proposed an optimization-based adversarial attack technique named CARROTA. CARROTA supports code transformations at both the token and statement levels and employs an iterative strategy to continuously search for adversarial examples that can mislead models into incorrect classifications. Unlike the MHM method, which only perturbs at the identifier level, CARROTA enables any grammar-compliant synonym transformations and integrates gradient information to guide the adversarial sample search process, thereby reducing computational overhead and improving attack efficiency. Experimental results demonstrate that CARROTA can efficiently generate adversarial examples, leading to an average performance drop of 87.2% at the token level, significantly outperforming MHM’s 75.5%. In addition, they also introduce a robustness evaluation metric, CARROTM, which approximates the robustness boundaries under stricter perturbation constraints, achieving significantly tighter robustness estimates compared to random baseline methods and MHM.
Zhang et al. (Zhang et al., 2023d) propose a attack method, STRUCK, targeting CodeLMs for code representation. STRUCK fully exploits the structural features of code to generate adversarial examples that mislead CodeLMs while maintaining the original semantics of the code. STRUCK integrates both global and local perturbation methods, using hierarchical structural features to select appropriate perturbations during adversarial example generation. Through comprehensive experiments on several code classification tasks, STRUCK demonstrates its superior effectiveness, efficiency, and imperceptibility compared to existing attacks like MHM and S-CARROTA (Zhang et al., 2022a). The experimental results show that STRUCK achieves a performance degradation of over 80% on various CodeLMs and significantly improves the robustness of models when used for adversarial training.
@@
To assess the robustness of software vulnerability detection models, Ferretti et al. (Ferretti and Saletta, 2021) propose an evolutionary approach. The fundamental idea of the method is to explore the potential feature representation space using the Grammatical Evolution (GE) algorithm (O’Neill et al., 2014). Initially, the method evolves code snippets by using the classifier’s output as the fitness function. Then, it examines how well the selected evolved individuals can effectively alter the classification of any instance in the labeled dataset. Finally, the method generates adversarial samples by injecting semantics while preserving the syntax. This method leverages evolutionary pressure rather than random sampling, thus effectively exploring the behavior of classifiers and identifying features that could potentially mislead the model. Their experiments indicate that code snippets generated by the GE algorithm can significantly guide the network to classify modified instances in a manner opposite to the true labels with high probability, reducing the precision/recall area under curve (P/R-AUC index) of the neural classifier from 0.42 to 0.06.